if tangent theta = square root of 2 over 4 and angle theta terminates in the first quadrant, determine co-tangent theta, co-secant theta, and secant theta
Answer by stanbon(75887)

if tangent theta = square root of 2 over 4 and angle theta terminates in the first quadrant, determine co-tangent theta, co-secant theta, and secant theta
-----
Tan is defined as y/x
So y = sqrt(2) and x = 4
Then r = sqrt[2+16] = sqrt(18)
---------
cot(t) = 4/sqrt(2)
csc(t) = r/y = sqrt(18)/sqrt(2) = sqrt(9) = 3
--------------------
sec(t) = r/x = sqrt(18)/4
===========================
Cheers,
Stan H.
---------------------------